# Curve fitting on a quantum annealer for an advanced navigation method

### v2

@@ IV.4 Further discussion
Recent literature has addressed this challenge using learning-based strategies, most notably through the approximation of the value function or the HJB operator via deep neural networks. Notable examples include the Deep Galerkin Method [68] and physics-informed neural networks [69]. Both utilize neural networks as function approximators trained on residuals of partial differential equations. These methods are well-suited for high-dimensional spaces by leveraging the expressive capacity and scalability of deep architectures.
In contrast to these continuous, gradient-based approaches, our method introduces a discrete, combinatorial framework for function approximation based on a QUBO. This discrete optimization perspective can be interpreted as a learning approach that bypasses gradient descent and instead exploits a new perspective on value function approximation, particularly in problem settings that can be nonconvex or nondifferentiable. In such cases, the resulting value function may exhibit sharp transitions or multiple local extrema. When approximating such functions using triangular functions on a discretized grid, the associated QUBO landscape can become irregular, potentially containing many local minima. This may lead to instabilities in the QUBO optimization and reduced approximation accuracy. However, this is not an intrinsic limitation of the QUBO approach itself, but rather a structural artifact of the discretization and basis choice. In fact, quantum annealing metaheuristics, such as those employed by D-Wave, offer the potential to escape local minima via quantum tunneling, which may be advantageous in overcoming the challenges of nonconvexity.
@@ V Summary and conclusions
Such approximations can often be reduced to curve fitting in one dimension. In this work, we formulated least-squares curve fitting, where we allowed for a rather general approximator in form of a finite linear combination of standardized (basis) functions, as a QUBO that is suited to be solved with quantum annealing. For simple functions/data points, we found that quantum annealing in its current state is in principle able to deliver comparable results to a classical computer. However, curve fitting tasks that result in a QUBO size larger than roughly about $60\times 60$ (i.e., with about 60 logical variables or more) cannot be solved accurate enough on a D-Wave annealer yet. The reason for that is the difficult low-energy landscape of the resulting Ising Hamiltonian. Moreover, QUBO problems described by global basis functions (e.g., Chebyshev polynomials) are even more difficult to solve, hence even smaller problems can have a poor solution quality.
@@
Nevertheless, the quantum annealing-oriented curve fitting principally works. Regarding a real-world use case—albeit still in an exploratory fashion—we investigated the linear-quadratic optimal control problem of just-in-time arrival of a vessel, where we approximated the value function in the Bellman equation. With an increasing number of basis functions, the solution space can be better represented, leading to more accurate value estimates compared to the grid-based method. Nevertheless, there was hope that the quantum annealing process would yield lower errors in larger search spaces. Unfortunately, we find that D-Wave’s current quantum annealing hardware does not yet appear to be capable of meeting this expectation.
Acknowledgments
We acknowledge funding and support from the German Federal Ministry for Economic Affairs and Climate Action (Bundesministerium für Wirtschaft und Klimaschutz) through the PlanQK initiative (01MK20005).
This project (HA project no. 1362/22-67) is financed with funds of LOEWE—Landes-Offensive zur Entwicklung Wissenschaftlich-ökonomischer Exzellenz, Förderlinie 3: KMU-Verbundvorhaben (State Offensive for the Development of Scientific and Economic Excellence).
Declarations
Author contributions
Conceptualization, funding acquisition, and supervision: B. Harrach and W. Mergenthaler. Formal analysis, investigation, and software: P. Isserstedt and D. Jaroszewski. Validation: P. Isserstedt, D. Jaroszewski, and F. Paul. Writing (original draft): P. Isserstedt and D. Jaroszewski. Writing (review and editing): P. Isserstedt, D. Jaroszewski, and F. Paul. This list uses the Contributor Role Taxonomy (CRediT) [71].
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.

### v3

@@ IV.4 Further discussion
Recent literature has addressed this challenge using learning-based strategies, most notably through the approximation of the value function or the HJB operator via deep neural networks. Notable examples include the Deep Galerkin Method [68] and physics-informed neural networks [69, 70]. Both utilize neural networks as function approximators trained on residuals of partial differential equations. These methods are well-suited for high-dimensional spaces by leveraging the expressive capacity and scalability of deep architectures.
In contrast to these continuous, gradient-based approaches, our method introduces a discrete, combinatorial framework for function approximation based on a QUBO. This discrete optimization perspective can be interpreted as a learning approach that bypasses gradient descent and instead exploits a new perspective on value function approximation, particularly in problem settings that can be nonconvex or nondifferentiable. In such cases, the resulting value function may exhibit sharp transitions or multiple local extrema. When approximating such functions using triangular functions on a discretized domain, the associated QUBO landscape can become irregular, potentially containing many local minima. This may lead to instabilities in the QUBO optimization and reduced approximation accuracy. However, this is not an intrinsic limitation of the QUBO formulation itself, but rather a structural artifact of the discretization and basis choice. In fact, quantum annealing metaheuristics, such as those employed by the D-Wave, offer the potential to escape local minima via quantum tunneling, which may be advantageous in overcoming the challenges of nonconvexity.
@@
Finally, we would like to note that the practical implementation of the speed profile optimization as detailed in Section˜IV.3 is based on a simplified scenario. This simplification is intentional, as the full optimization problem rapidly becomes computationally expensive due to the curse of dimensionality. Therefore, we deliberately considered a reduced, one-dimensional model to illustrate the dynamic programming approach. Despite its simplicity, our model already incorporates an important factor, the magnitude of the projected sea current, which directly affects the optimal speed profile and fuel consumption. Moreover, the proposed framework is flexible and can be readily extended to include additional physical effects. For example, it could be extended to account for wind and wave conditions, variations depending on ship draft or trim, or even to compute an optimal acceleration profile instead of a speed-only control. These extensions would allow a more realistic representation of maritime operations while maintaining the systematic structure provided by the dynamic programming approach. Such a comprehensive study, however, is beyond the scope of the present work and is left for future research.
@@ V Summary and conclusions
Such approximations can often be reduced to curve fitting in one dimension. In this work, we formulated least-squares curve fitting, where we allowed for a rather general approximator in form of a finite linear combination of standardized (basis) functions, as a QUBO that is suited to be solved with quantum annealing. For simple functions/data points, we found that quantum annealing in its current state is in principle able to deliver comparable results to a classical computer. However, curve fitting tasks that result in a QUBO size larger than roughly about $60\times 60$ (i.e., with about 60 logical variables or more) cannot be solved accurate enough on a D-Wave annealer yet. The reason for that is the difficult low-energy landscape of the resulting Ising Hamiltonian. Moreover, QUBO problems described by global basis functions (e.g., Chebyshev polynomials) are even more difficult to solve, such that even smaller problems can have poor quality.
@@
Nevertheless, the quantum annealing-oriented curve fitting principally works. Regarding a real-world use case—albeit still in an exploratory fashion—we investigated the linear-quadratic optimal control problem of just-in-time arrival of a vessel, where we approximated the value function in the Bellman equation. With an increasing number of basis functions, the solution space can be better represented, leading to more accurate value estimates compared to the grid-based method. There was hope that the quantum annealing process would yield lower errors in larger search spaces. Unfortunately, we find that D-Wave’s current quantum annealing hardware does not yet appear to be capable of meeting this expectation.
Finally, we note that the present work focuses on analyzing the accuracy and solution quality of the quantum annealing approach to curve fitting and speed profile optimization under well-defined and controlled conditions. For this reason, potential performance enhancements through QUBO preprocessing techniques (e.g., sparsification) and comparisons with quantum-inspired classical solvers like SimCIM or NGQ are not considered at this stage. Runtime optimization and scaling behavior represent important directions for future research. Such investigations will help to clarify whether the current limitations of quantum annealing are intrinsic to the method itself or primarily related to currently existing hardware constraints. By first establishing a reliable accuracy baseline, this study lays the groundwork for us for subsequent studies on quantum performance, efficiency, and practical applicability.
Acknowledgments.   We acknowledge funding and support from the German Federal Ministry for Economic Affairs and Climate Action (Bundesministerium für Wirtschaft und Klimaschutz) through the PlanQK initiative (01MK20005). This project (HA project no. 1362/22-67) is financed with funds of LOEWE—Landes-Offensive zur Entwicklung Wissenschaftlich-ökonomischer Exzellenz, Förderlinie 3: KMU-Verbundvorhaben (State Offensive for the Development of Scientific and Economic Excellence).
Author contributions.   Conceptualization, funding acquisition, and supervision: WM, BH. Formal analysis, investigation, and software: PI, DJ. Validation: PI, DJ, FP. Writing (original draft): PI, DJ. Writing (review and editing): PI, DJ, FP.
Competing interests.   The authors have no competing interests to declare that are relevant to this article.